OpenStudy (anonymous):

This prism has a volume of 112 cm3. What would the volume of the prism be if each dimension was doubled? (Scale factor is 2.)

OpenStudy (anonymous):

i would help, but someone already answered the same question a long time ago here's the link to the question, and answer: http://openstudy.com/updates/51343c34e4b0034bc1d80b83

OpenStudy (jdoe0001):

\(\bf volume=length\cdot width\cdot height \\ \quad \\ \textit{doubling all 3 sides}\implies volume=2length\cdot 2width\cdot 2height\)

OpenStudy (jdoe0001):

you could see it this way v = lwh if we double l and w and h then v = 2l2w2h v= 2*2*2lwh v = 8lwh v = 8(lwh)

OpenStudy (anonymous):

This prism has a volume of 112 cm3. What would the volume of the prism be if each dimension was doubled? (Scale factor is 2.) A. 896 cm3 B. 224 cm3 C. 336 cm3 D. 448 cm3

OpenStudy (anonymous):

I am confused